# Super-Linear: A Lightweight Pretrained Mixture of Linear Experts for Time Series Forecasting

### v1

@@ Abstract
Time series forecasting (TSF) is critical in domains like energy, finance, healthcare, and logistics, requiring models that generalize across diverse datasets. Large pre-trained models such as Chronos and Time-MoE show strong zero-shot (ZS) performance but suffer from high computational costs. In this work, We introduce Super-Linear, a lightweight and scalable mixture-of-experts (MoE) model for general forecasting. It replaces deep architectures with simple frequency-specialized linear experts, trained on resampled data across multiple frequency regimes. A lightweight spectral gating mechanism dynamically selects relevant experts, enabling efficient, accurate forecasting. Despite its simplicity, Super-Linear matches state-of-the-art performance while offering superior efficiency, robustness to various sampling rates, and enhanced interpretability. The implementation of Super-Linear is available at https://github.com/azencot-group/SuperLinear
Machine Learning, ICML
@@ 1 Introduction
Time series forecasting plays a vital role in many high-impact domains, from optimizing energy grids and managing financial risk to supporting clinical decision-making and supply chain operations (Taylor & Letham, 2018; Salinas et al., 2020; Lim et al., 2021). As these systems increasingly rely on automated forecasting tools, the demand grows for models that can generalize across a wide range of temporal settings without manual tuning. However, many real-world scenarios involve small or heterogeneous datasets with unique periodic patterns, limited labels, and shifting dynamics (Norton et al., 2025). In such environments, training specialized models per dataset is computationally costly and often impractical. This motivates the need for zero-shot forecasting (ZSF), a setting in which a single model can generalize to entirely unseen datasets and domains without retraining or fine-tuning (Shi et al., 2024; Liu et al., 2024d, c; Das et al., 2024).
Figure 1: Performance versus inference time trade-off across different prominent pretrained TSF models.
@@
Recent work has made significant progress toward this goal. Large pretrained models such as Timer-XL, TimeMoE, and TimesFM (Liu et al., 2024c; Shi et al., 2024; Das et al., 2024) employ transformer-based backbones, input patching, and masked pretraining to achieve strong zero-shot performance across domains. However, these gains come at a high computational cost. For instance, TimeMoE was trained over several days using 28 A100-80GB GPUs, whereas Chronos can take over 10 seconds to generate an output for a batch size of 64 under certain configurations. See Fig. 1 for a general overview. Additionaly, the complexity of these architectures also poses a challenge for interpretability, as their sophisticated designs make it difficult to understand or explain forecasting decisions. Developing models that combine strong generalization with greater transparency still remains a key challenge, which often hinders accessibility and practical deployment. As zero-shot forecasting becomes more practical, it is imperative to develop models that preserve strong generalization while dramatically reducing the computational and memory footprint.
@@
Recent findings show that carefully constructed linear models can outperform deep transformer-based architectures across several TSF benchmarks (Zeng et al., 2023; Toner & Darlow, 2024; Ni et al., 2024; Nochumsohn et al., 202t). Consequently, linear forecasting models offer a promising foundation for efficient ZSF and full-shot forecasting (FSF). They are fast, lightweight, and well-aligned with the temporal structure of many natural sequences. However, a single linear model typically lacks the expressivity needed to generalize across diverse datasets (Nochumsohn et al., 202t). Capturing multiple overlapping frequencies with one parameter set induces what we term frequency confusion, where different temporal patterns compete for representation and degrade accuracy. This limitation is especially pronounced in zero-shot settings, where effective generalization requires both robustness to distribution shifts and adaptability to novel temporal patterns. Taken together, these challenges may explain why linear models remain relatively underexplored in zero-shot forecasting within pre-trained frameworks.
@@ 2 Related Work
Recent advancements in pretrained models for vision and text (Zhou et al., 2024) have accelerated the development of large time series foundation models for zero-shot, few-shot, and full-shot forecasting (Liang et al., 2024). Modern models like UniTime (Liu et al., 2024b), TEMPO (Cao et al., 2023), and GPT4TS (Zhou et al., 2023) adopted GPT-based architectures, inspiring works such as MOMENT (Goswami et al., 2024) and TimesFM (Das et al., 2024), which integrate transformers (Vaswani et al., 2017), input patching, and masked training to handle diverse inputs. Chronos (Ansari et al., 2024) leverages a T5-based architecture (Raffel et al., 2020) with discrete time tokenization, while Moirai (Woo et al., 2024) employs an encoder-only transformer with multi-patch size projections and binary attention biases for universal forecasting. Timer and Timer-XL (Liu et al., 2024d, c), as decoder-only generative models, use next-token prediction for flexible time series processing. Finally, Time-MoE (Shi et al., 2024), a decoder-only mixture-of-experts model, effectively scales parameters, enhancing accuracy and efficiency in forecasting.
@@
Before deep time series forecasting (TSF) gained traction, statistical and linear models such as ARIMA, Exponential Smoothing, and least squares regression (Shumway & Stoffer, 2000) dominated TSF. In recent years, MLP (multilayer perceptron) and transformer-based models initially outperformed these methods with models such as Informer, Autoformer, FEDformer, and N-HiTS (Zhou et al., 2021; Wu et al., 2021; Zhou et al., 2022; Challu et al., 2023), however, recent research has demonstrated that linear and sparse models such as DLinear and RLinear (Zeng et al., 2023; Toner & Darlow, 2024; Li et al., 2023) can be competitive or even superior. This has sparked interest in sparse and frequency-aware TSF models, including FITS and SparseTSF (Xu et al., 2023; Lin et al., 2024b), which leverage the frequency domain for improved forecasting. MTLinear (Nochumsohn et al., 202t) addresses gradient conflicts with variate cross-correlation, enhancing robustness against transformers with linear models, while CycleNet (Lin et al., 2024a) employs a shallow MLP with learnable recurrent cycles to model periodic patterns efficiently. These approaches provide scalable and interpretable forecasting solutions, particularly for resource-constrained applications.
@@
Mixture of Experts (MoE) (Cai et al., 2024; Jacobs et al., 1991) enhances efficiency and scalability by dynamically selecting specialized sub-models for different tasks, reducing computational overhead while maintaining high performance. Although widely used in language models (Cai et al., 2024), MoE has been sparingly applied to time series forecasting, beginning with the Mixture of Experts Model (MEM) (Zeevi et al., 1996), which extends autoregressive models into a nonlinear universal approximator. Models such as FEDformer (Zhou et al., 2022) utilize MoE for different average pooling selection, while Moirai-MoE and Time-MoE (Woo et al., 2024; Shi et al., 2024), incorporate MoE within their designs. MoLE (Ni et al., 2024) introduces weighted linear experts for in-domain forecasting, and FreqMoE (Liu, 2025) processes distinct frequency bands for prediction. Our approach scales these ideas into a centralized, linear-enhanced pretrained model capable of handling diverse datasets while maintaining a sparse architectural design.
@@ 3 Super-Linear Mixture-of-Experts
Given a historical sequence of observations $X_{1:L}=[x_{1},x_{2},\dots,x_{L}]^{T}\in\mathbb{R}^{L}$, where $L$ represents the number of lookback time steps, the goal in time series forecasting is to predict the future $H$ steps $Y_{1:H}=X_{L+1:L+H}=[x_{L+1},\dots,x_{L+H}]^{T}\in\mathbb{R}^{H}$. Super-Linear MoE leverages a modular architecture combining linear maps and expert gating mechanisms to enhance forecasting accuracy across diverse horizons with a single model, ranging from short-term (e.g., one step) to long-term (e.g., 720 steps and beyond). By dynamically selecting relevant experts based on input characteristics, the model adapts to varying frequencies and amplitudes. Furthermore, the integration of channel-wise independence ensures scalability to high-dimensional datasets, enabling Super-Linear to generalize effectively across multivariate forecasting tasks in real-world scenarios.
@@
Time series is naturally periodic, yet common methods, such as seasonal-trend decomposition or covariate timestamp embeddings (Ni et al., 2024), struggle with scalability and mixed-frequency signals (Theodosiou, 2011; Woo et al., 2022; Qin et al., 2024). Inspired by Fourier analysis (Shumway & Stoffer, 2000), we view a time series as a weighted sum of its dominant periodic components, with $\hat{\omega}_{i}$ denoting the characteristic frequencies and $\alpha_{i}$ are their amplitudes,
@@
Linear forecasting models that take the form of an autoregressive process remain surprisingly effective (Zeng et al., 2023), and often learn to emphasize seasonal lags—e.g., daily or weekly offsets—when such structure exists. However, capturing multiple overlapping frequencies with a single weight vector is fundamentally limited. As different seasonalities compete, shared weights induce “frequency confusion” and degrade performance (see Fig. 2). To address this, we propose approximating forecasts using a mixture of frequency-specialized linear experts. Given input $X\in\mathbb{R}^{B\times T}$ and output $Y\in\mathbb{R}^{B\times H}$, where $B$ is the batch size and $T$ and $H$ are the lookback and horizon, we have:
@@
Let $X\in\mathbb{R}^{B\times T}$ and $Y\in\mathbb{R}^{B\times H}$. Super-Linear produces a forecast $\hat{Y}\in\mathbb{R}^{B\times H}$ over a horizon of $H$ steps. This forecast is computed as a weighted sum over $N$ experts, each of which contributes a prediction $F_{i}(X)\in\mathbb{R}^{B\times H}$, modulated by a gating weight $G_{i}(X)\in\mathbb{R}^{B}$:
$\hat{Y}=\sum_{i=1}^{N}\text{diag}(G_{i}(X))F_{i}(X)\ ,$ (3)
where $\mathrm{diag}(A)$ is the diagonal matrix in $\mathbb{R}^{B\times B}$ with the entries of $A\in\mathbb{R}^{B}$ on its main diagonal.
Figure 3: Super-Linear architecture overview. A frequency-aware gating router computes sparse scores from the input frequencies, dynamically selecting a subset of linear experts (1) including pre-trained frequency specialists, complementary modules, and heuristic naïve and mean baselines (2) whose predictions are combined to produce the final forecast (3).
@@
Each expert $F_{i}$ is a learnable or non-learnable linear model. The learnable experts consist of $N_{f}$ frequency-specific modules and $N_{c}$ complementary ones, implemented via RLinear—a linear layer combined with instance normalization using $\text{RevIN}_{\text{norm}}$ (Toner & Darlow, 2024). This normalization mitigates distribution shifts by reapplying the input’s original scale after transformation. The output of each learnable expert is:
$F_{i}(X)=\text{RevIN}^{-1}\left(\text{RevIN}(X)W_{i}+\text{rep}(b_{i})\right)\ ,$ (4)
where $W_{i}\in\mathbb{R}^{T\times H}$ and $b_{i}\in\mathbb{R}^{H}$ are trainable, and $\text{rep}(\cdot)$ repeats its input $B$ times. Frequency experts are pre-trained for a specific $\hat{\omega}_{i}$ (App. C.3), covering diverse periodicities, while complementary experts are trained jointly to capture residual temporal signals. Two non-learnable experts, the naïve (last value) and mean (input average), serve as robust heuristics in low-frequency or short-sequence regimes. Including them improves performance across benchmarks (Tab. 3).
@@
The weights $G_{i}(X)$ for combining expert outputs are derived through a sparse gating mechanism (Shazeer et al., 2017) that dynamically selects the most relevant experts based on the input’s spectral properties. Specifically, we compute a gating matrix $g(X)\in\mathbb{R}^{B\times N}$, where each entry reflects the affinity between the input and a given expert. These entries are computed using the normalized periodogram (Shumway & Stoffer, 2000; Schuster, 1898) of the zero-mean input. Formally,
@@
To encourage sparsity and specialization, only the top-$k$ scoring experts are activated for each input. During training, we inject Gaussian noise $\mathcal{R}^{\text{noise}}\in\mathbb{R}^{B\times N}$ into the scores to promote exploration. The final gating weights, $G(X)\in\mathbb{R}^{B\times N}$, are computed using a softmax applied only over the top-$k$ scores per column, i.e., TopK eliminates the $B\times(N-k)$ smallest elements of $g(X)+\sigma\mathcal{R}^{\text{noise}}$ for $\sigma=0.1$:
@@
Super-Linear adopts a two-stage training framework, illustrated in Fig. 4: Stage 1: Frequency expert pre-training. Each expert is trained independently on data aligned with its designated fundamental frequency $\hat{\omega}_{i}$. To expose each expert to diverse temporal patterns, we apply extensive resampling augmentations using linear interpolation. Unlike prior work focused on dataset diversity, our emphasis is on frequency abundance, capturing a broad range of frequency instances rather than dataset heterogeneity. We thus select a subset of datasets from Monash (Godahewa et al., 2021) and PEMS (Liu et al., 2022) that provide sufficient sequence length for downsampling. Full details on dataset selection and resampling are provided in App. B.

### v2

@@ Abstract
Time series forecasting (TSF) is critical in domains like energy, finance, healthcare, and logistics, requiring models that generalize across diverse datasets. Large pre-trained models such as Chronos and Time-MoE show strong zero-shot (ZS) performance but suffer from high computational costs. In this work, we introduce Super-Linear, a lightweight and scalable mixture-of-experts (MoE) model for general forecasting. It replaces deep architectures with simple frequency-specialized linear experts, trained on resampled data across multiple frequency regimes. A lightweight spectral gating mechanism dynamically selects relevant experts, enabling efficient, accurate forecasting. Despite its simplicity, Super-Linear demonstrates strong performance across benchmarks, while substantially improving efficiency, robustness to sampling rates, and interpretability. The implementation of Super-Linear is available at: https://github.com/azencot-group/SuperLinear.
@@ 1 Introduction
Time series forecasting plays a vital role in many high-impact domains, from optimizing energy grids and managing financial risk to supporting clinical decision-making and supply chain operations Taylor and Letham (2018); Salinas et al. (2020); Lim et al. (2021). As these systems increasingly rely on automated forecasting tools, the demand grows for models that can generalize across a wide range of temporal settings without manual tuning. However, many real-world scenarios involve small or heterogeneous datasets with unique periodic patterns, limited labels, and shifting dynamics Norton et al. (2025). In such environments, training specialized models per dataset is computationally costly and often impractical. This motivates the need for zero-shot forecasting (ZSF), a setting in which a single model can generalize to entirely unseen datasets and domains without retraining or fine-tuning Shi et al. (2024); Liu et al. (2024d; c); Das et al. (2024).
@@
Recent work has made significant progress toward this goal. Large pretrained models such as Timer-XL, TimeMoE, and TimesFM (Liu et al., 2024c; Shi et al., 2024; Das et al., 2024) employ transformer-based backbones, input patching, and masked pretraining to achieve strong zero-shot performance across domains. However, these gains come at a high computational cost. For instance, TimeMoE was trained over several days using 28 A100-80GB GPUs, whereas Chronos can take over 4 seconds to generate an output for a batch size of 64 under certain configurations. See Fig. 1 for a general overview. Additionally, the complexity of these architectures also poses a challenge for interpretability, as their sophisticated designs make it difficult to understand or explain forecasting decisions. Developing models that combine strong generalization with greater transparency still remains a key challenge, which often hinders accessibility and practical deployment. As zero-shot forecasting becomes more practical, it is imperative to develop models that preserve strong generalization while dramatically reducing the computational and memory footprint.
Figure 1: Performance versus inference time trade-off across different prominent pretrained TSFM on the GIFT-Eval and LTSF benchmarks.
@@
Recent findings show that carefully constructed linear models can outperform deep transformer-based architectures across several TSF benchmarks (Zeng et al., 2023; Toner and Darlow, 2024; Ni et al., 2024; Nochumsohn et al., 202t). Consequently, linear forecasting models offer a promising foundation for efficient ZSF and full-shot forecasting (FSF). They are fast, lightweight, and well-aligned with the temporal structure of many natural sequences. However, a single linear model typically lacks the expressivity needed to generalize across diverse datasets (Nochumsohn et al., 202t). Capturing multiple overlapping frequencies with one parameter set induces what we term frequency confusion, where different temporal patterns compete for representation and degrade accuracy. This limitation is especially pronounced in zero-shot settings, where effective generalization requires both robustness to distribution shifts and adaptability to novel temporal patterns. Taken together, these challenges may explain why linear models remain relatively underexplored in zero-shot forecasting within pre-trained frameworks.
@@ 2 Related Work
Recent advancements in pretrained models for vision and text (Zhou et al., 2024) have accelerated the development of large time series foundation models for zero-shot, few-shot, and full-shot forecasting (Liang et al., 2024). Modern models like UniTime (Liu et al., 2024b), TEMPO (Cao et al., 2023), and GPT4TS (Zhou et al., 2023) adopted GPT-based architectures, inspiring works such as MOMENT (Goswami et al., 2024) and TimesFM (Das et al., 2024), which integrate transformers (Vaswani et al., 2017), input patching, and masked training to handle diverse inputs. Chronos (Ansari et al., 2024) leverages a T5-based architecture (Raffel et al., 2020) with discrete time tokenization, while Moirai (Woo et al., 2024) employs an encoder-only transformer with multi-patch size projections and binary attention biases for universal forecasting. Timer and Timer-XL (Liu et al., 2024d; c), as decoder-only generative models, use next-token prediction for flexible time series processing. Finally, Time-MoE (Shi et al., 2024), a decoder-only mixture-of-experts model, effectively scales parameters, enhancing accuracy and efficiency in forecasting.
@@
Before deep time series forecasting (TSF) gained traction, statistical and linear models such as ARIMA, Exponential Smoothing, and least squares regression (Shumway and Stoffer, 2000) dominated TSF. In recent years, MLP (multilayer perceptron) and transformer-based models initially outperformed these methods with models such as Informer, Autoformer, FEDformer, and N-HiTS (Zhou et al., 2021; Wu et al., 2021; Zhou et al., 2022; Challu et al., 2023), however, recent research has demonstrated that linear and sparse models such as DLinear and RLinear (Zeng et al., 2023; Toner and Darlow, 2024; Li et al., 2023) can be competitive or even superior. This has sparked interest in sparse and frequency-aware TSF models, including FITS and SparseTSF (Xu et al., 2023; Lin et al., 2024b), which leverage the frequency domain for improved forecasting. MTLinear (Nochumsohn et al., 202t) addresses gradient conflicts with variate cross-correlation, enhancing robustness against transformers with linear models, while CycleNet (Lin et al., 2024a) employs a shallow MLP with learnable recurrent cycles to model periodic patterns efficiently. These approaches provide scalable and interpretable forecasting solutions, particularly for resource-constrained applications.
@@
Mixture of Experts (MoE) (Cai et al., 2024; Jacobs et al., 1991) enhances efficiency and scalability by dynamically selecting specialized sub-models for different tasks, reducing computational overhead while maintaining high performance. Although widely used in language models (Cai et al., 2024), MoE has been sparingly applied to time series forecasting, beginning with the Mixture of Experts Model (MEM) (Zeevi et al., 1996), which extends autoregressive models into a nonlinear universal approximator. Models such as FEDformer (Zhou et al., 2022) utilize MoE for different average pooling selection, while Moirai-MoE and Time-MoE (Woo et al., 2024; Shi et al., 2024), incorporate MoE within their designs. MoLE (Ni et al., 2024) introduces weighted linear experts for in-domain forecasting, and FreqMoE (Liu, 2025) processes distinct frequency bands for prediction. Distinguishably, our approach scales these ideas into a centralized, linear-enhanced pretrained model capable of handling diverse datasets, while maintaining a sparse architecture built around a single linear-expert MoE layer, unlike other MoE methods that rely on intermediate, non-linear modules.
@@ 3 Super-Linear Mixture-of-Experts
Given a historical sequence of observations $X_{1:L}=[x_{1},x_{2},\dots,x_{L}]^{T}\in\mathbb{R}^{L}$, where $L$ represents the number of lookback time steps, the goal in time series forecasting is to predict the future $H$ steps $Y_{1:H}=X_{L+1:L+H}=[x_{L+1},\dots,x_{L+H}]^{T}\in\mathbb{R}^{H}$. Super-Linear MoE leverages a modular architecture combining linear maps and expert gating mechanisms to enhance forecasting accuracy across diverse horizons with a single model, ranging from short-term (e.g., one step) to long-term (e.g., 720 steps and beyond). By dynamically selecting relevant experts based on input characteristics, the model adapts to varying frequencies and amplitudes. Furthermore, the model design employs channel-wise independence (univariate) to ensure scalability to high-dimensional datasets, enabling Super-Linear to generalize effectively across multivariate forecasting tasks in real-world scenarios.
In this work, we evaluate Super-Linear under both zero-shot (ZS) and in-domain (full-shot, FS) forecasting settings. In our usage, zero-shot denotes applying the pretrained model to entirely unseen datasets with no additional training or adaptation, whereas in-domain forecasting refers to performance after training the model directly on the target dataset.
@@
Time series is naturally periodic, yet common methods, such as seasonal-trend decomposition or covariate timestamp embeddings (Ni et al., 2024), struggle with scalability and mixed-frequency signals (Theodosiou, 2011; Woo et al., 2022; Qin et al., 2024). Inspired by Fourier analysis (Shumway and Stoffer, 2000), we view a time series as a weighted sum of its dominant periodic components, with $\hat{\omega}_{i}$ denoting the characteristic frequencies and $\alpha_{i}$ are their amplitudes,
@@
Linear forecasting models that take the form of an autoregressive process remain surprisingly effective Zeng et al. (2023), and often learn to emphasize seasonal lags—e.g., daily or weekly offsets—when such structure exists. However, capturing multiple overlapping frequencies with a single weight vector is fundamentally limited. As different seasonalities compete, shared weights induce “frequency confusion” and degrade performance (see Fig. 2). To address this, we propose approximating forecasts using a mixture of frequency-specialized linear experts. Given input $X\in\mathbb{R}^{1\times T}$ and output $Y\in\mathbb{R}^{1\times H}$, where $T$ and $H$ are the lookback and horizon, we have:
@@
Let $X\in\mathbb{R}^{1\times T}$ and $Y\in\mathbb{R}^{1\times H}$. Super-Linear produces a forecast $\hat{Y}\in\mathbb{R}^{1\times H}$ over a horizon of $H$ steps. This forecast is computed as a weighted sum over $N$ experts, each of which contributes a prediction $F_{i}(X)\in\mathbb{R}^{1\times H}$, modulated by a gating weight $G_{i}(X)\in\mathbb{R}$:
$\hat{Y}=\sum_{i=1}^{N}G_{i}(X)F_{i}(X)\ ,$ (3)
Figure 3: Super-Linear architecture overview. A frequency-aware gating router computes sparse scores from the input frequencies, dynamically selecting a subset of linear experts, (1) including linear experts (2) whose predictions are combined to produce the final forecast (3).
@@
Each expert $F_{i}$ is a learnable or non-learnable linear model. The learnable experts consist of $N_{f}$ frequency-specific modules and $N_{c}$ complementary ones, implemented via RLinear—a linear layer combined with instance normalization using RevIN Toner and Darlow (2024). This normalization mitigates distribution shifts by reapplying the input’s original scale after transformation. The output of each learnable expert is:
$F_{i}(X)=\text{RevIN}^{-1}\left(\text{RevIN}(X)W_{i}+b_{i}\right)\ ,$ (4)
where $W_{i}\in\mathbb{R}^{T\times H}$ and $b_{i}\in\mathbb{R}^{H}$ are trainable. Frequency experts are pre-trained with data associated with a specific frequency $\hat{\omega}_{i}$, covering diverse periodicities, while complementary experts are trained jointly to capture residual temporal signals. More on the frequencies, and how they were obtained, is given in App. D.3). Two non-learnable experts, the naïve (last value) and mean (input average), serve as robust heuristics in low-frequency or short-sequence regimes. Including them improves performance across benchmarks (Tab. 2).
@@
The weights $G_{i}(X)$ for combining expert outputs are derived through a sparse gating mechanism (Shazeer et al., 2017) that dynamically selects the most relevant experts based on the input’s spectral properties. Specifically, we compute a gating matrix $g(X)\in\mathbb{R}^{1\times N}$, where each entry reflects the affinity between the input and a given expert. These entries are computed using the normalized periodogram (Shumway and Stoffer, 2000; Schuster, 1898) of the zero-mean input. Formally,
@@
To encourage sparsity and specialization, only the top-$k$ scoring experts are activated for each input. During training, we inject Gaussian noise $\mathcal{R}^{\text{noise}}\in\mathbb{R}^{1\times N}$ into the scores to promote exploration. The final gating weights, $G(X)\in\mathbb{R}^{1\times N}$, are computed using a softmax applied only over the top-$k$ scores per column, i.e., TopK eliminates the (Nk)(N-k) smallest elements of $g(X)+\sigma\mathcal{R}^{\text{noise}}$ for $\sigma=0.1$:
@@
Super-Linear adopts a two-stage training framework, illustrated in Fig. 4: Stage 1: Frequency expert pre-training. Each expert is trained independently on data aligned with its designated fundamental frequency $\hat{\omega}_{i}$. To expose each expert to diverse temporal patterns, we apply extensive resampling augmentations using linear interpolation. Unlike prior work focused on dataset diversity, our emphasis is on frequency abundance, capturing a broad range of frequency instances rather than dataset heterogeneity. We thus select a subset of datasets from Monash Godahewa et al. (2021) and PEMS Liu et al. (2022) that provide sufficient sequence length for downsampling. Full details on dataset selection and resampling are provided in App. C.
@@
4 Experiments
4.1 Experimental Details
Pre-training Datasets and Resampling.
This work emphasizes frequency diversity over sheer dataset volume. For pre-training, we use two widely adopted sources: Monash (Godahewa et al., 2021) and PEMS (Liu et al., 2022). From Monash, only datasets with more than 5,000 samples are included, ensuring sufficient data for resampling. Resampling, both upsampling and downsampling, is performed on-the-fly during training via linear interpolation, eliminating the need to store multiple copies and saving space. We use the frequencies detailed in App. D.3. The same pre-training data is used across both training stages, as outlined in Sec. 3.
Training and Implementation.
All Super-Linear models are trained on a single NVIDIA RTX3090 (24GB) using PyTorch (Paszke et al., 2019). In Stage 1, linear experts are trained independently and later combined into the Super-Linear architecture. We use the Adam optimizer with MSE loss and without auxiliary loss factors. For full-shot, the top-k experts are chosen via validation loss. All Super-Linear models use a lookback window of 512 and forecast a horizon of 96; longer horizons are predicted autoregressively. Full hyperparameters are detailed in App. F.
Figure 5: GIFT-Eval performance and parameter count of Super-Linear compared to prominent foundation models in TSF . The MASE score represents the geometric mean MASE across datasets, normalized by the seasonal-naive.
Baselines and Evaluations.
We compare our method to top-performing models. For zero-shot, we benchmark against Timer-XL, ROSE, Time-MoE, Moirai, TimesFM variants, Chronos variants, Sundial, and TiRex (Liu et al., 2024c; Wang et al., ; Shi et al., 2024; Woo et al., 2024; Das et al., 2024; Ansari et al., 2024; Liu et al., 2025; Auer et al., 2025), and focus mostly on models available on GitHub or Hugging Face. Evaluation follows the LTSF benchmark in both zero-shot and full-shot setups, using ETTm1, ETTm2, ETTh1, ETTh2, Electricity, Traffic, and Weather datasets Liu et al. (2024c); Shi et al. (2024), including the GIFT-Eval benchmark (Aksu et al., 2024), which covers a wide range of dataset domains and sampling rates. The appendix includes additional evaluations such as full-shot (see Tab. 4), and extra model comparisons.
4.2 Main Results